OpenStudy (anonymous):

y^{'}=sqrt{\frac{ 1-y^{2} }{ 1-x ^{2} }}

OpenStudy (unklerhaukus):

\[y^{'}=\sqrt{\frac{ 1-y^{2} }{ 1-x ^{2} }}\]

OpenStudy (anonymous):

yes...i hve open the square root to equal 1/2

OpenStudy (dumbcow):

separate variables \[\frac{dy}{\sqrt{1-y^{2}}} = \frac{dx}{\sqrt{1-x^{2}}}\] integrate \[\sin^{-1} y = \sin^{-1} x +C\] \[y = \sin (\sin^{-1} x + C)\]

OpenStudy (anonymous):

can explain more how u can get sin\[\sin^{-1} x\]

OpenStudy (anonymous):

it use subtitude u

OpenStudy (dumbcow):

sure use trig substitution \[x = \sin \theta\] \[dx = \cos \theta\] \[\int\limits \frac{dx}{\sqrt{1-x^{2}}} = \int\limits \frac{\cos \theta}{\sqrt{1-\sin^{2} \theta}} = \int\limits d \theta\]

OpenStudy (anonymous):

still not understand...its the equation t \[\int\limits d \theta = \sin ^{-1} x\]

OpenStudy (dumbcow):

\[\int\limits d \theta = \theta\] we know \[x = \sin \theta\] to solve for theta in terms of "x" you must use inverse sine function \[\theta = \sin^{-1} x\]

OpenStudy (anonymous):

ohh ok undertsnd,,thanks for helping,,

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

its \[\sqrt{1-\sin ^{2} \theta}= \cos \theta \] from trig identity.. so we just simplify get just \[\int\limits d \theta\]

OpenStudy (anonymous):

to confirm the answer it true or not...and to more undertsand..TQQQ

OpenStudy (dumbcow):

oh yes correct...sorry i skipped some steps in my explanation from pythagorean thm \[\sin^{2} x + \cos^{2} x = 1\] we know \[\cos x = \sqrt{1-\sin^{2} x}\] then i simplified \[\frac{\cos \theta}{\cos \theta} d \theta = d \theta\]

OpenStudy (anonymous):

it okay u are lot helping just right know...hopely if anything i cannot understand i can ask u again???

OpenStudy (dumbcow):

sure if im on and available i'll help you out